# Limit Theorems and Quantitative Statistical Stability for the Equilibrium States of Piecewise Partially Hyperbolic Maps

### v1

@@ 5.2. Exponential Decay of Correlations over $L^{1}(\mathcal{F}_{0})$
In this section we introduce the space $L^{1}(\mathcal{F}_{0})$, on which we will later establish exponential decay of correlations. This result will have a further application, where we prove the Central Limit Theorem for Hölder observables.
@@
Theorem B.
Suppose that $F:\Sigma\longrightarrow\Sigma$ satisfies (f1), (f2), (f3), (H1) and (H2) and $(\alpha\cdot L)^{\zeta}<1$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi\in\mathcal{P}$ in $S^{\infty}$. There exists a constant $0<\tau_{3}<1$ such that, for all $\psi\in L^{1}_{\mu_{0}}(\mathcal{F}_{0})$ and all $\varphi\in\operatorname{H}_{\zeta}(\Sigma)$, it holds
@@
Note that if $\psi$ is a real-valued $\mathcal{F}_{0}$-measurable function, then $\psi$ is constant along the fibers. Indeed, $\psi^{-1}(x)$ is $\mathcal{F}_{0}$-measurable. If $\psi^{-1}(x)\neq\emptyset$, then there exist $\gamma\in\mathcal{F}^{s}$ and $z\in\gamma$ such that $\psi(z)=x$. By Proposition 5.3, we have $\gamma\subset\psi^{-1}(x)$, and therefore $\psi(z)=x$ for all $z\in\gamma$.
@@
By Theorem E, the proof is finished.
@@ 5.3. The Gordin’s Theorem
We now introduce an important tool that will be used, together with Theorem F, to establish the Central Limit Theorem: Gordin’s theorem. Its proof can be found in [21], and therefore we omit it here.
Let $(\Sigma,\mathcal{F}_{0},\mu)$ be a probability space where $\mathcal{F}_{0}$ is a $\sigma$-algebra of $\Sigma$. Let $F:\Sigma\longrightarrow\Sigma$ be a $\mathcal{F}_{0}$-measurable function and let $\mathcal{F}_{n}:=F^{-n}(\mathcal{F}_{0})$ be a nonincreasing family of $\sigma$-algebras. A function $\xi:M\to\mathbb{R}$ is $\mathcal{F}_{n}$-measurable if and only if there exists a $\mathcal{F}$-measurable function $\xi_{n}$ such that $\xi=\xi_{n}\circ F^{n}$. Define
@@ 5.4. Application: Central Limit Theorem
By the Theorem F, if $\psi\in L^{1}_{\mu_{0}}(\mathcal{F}_{0})$ with $\int\psi\,d\mu\leq 1$, then
@@
Theorem C (Central Limit Theorem).
Suppose that $F:\Sigma\longrightarrow\Sigma$ satisfies (f1), (f2), (f3), (H1) and (H2) and $(\alpha\cdot L)^{\zeta}<1$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi\in\mathcal{P}$ in $S^{\infty}$. Given a Hölder continuous function $\varphi$, define
@@ 6.1. Admissible $R(\delta)$-Perturbations
Since $\phi_{1}=1$ $m$-a.e., $\mu|_{\gamma_{\delta,i}}-\mu|_{\gamma}$ has zero average. Therefore, by lemmas 4.9 and 4.10, by (U2.2) and Definition (3.6) applied on $\mu$, we get
@@
for $m$-a.e. $\gamma\in M$ and all $\mu\in\mathcal{AB}_{m}$, where
@@
Understanding the action of the transfer operators $\{\mathop{\mathrm{\overline{}}{F}}_{\Phi,\delta}\}_{[0,\delta_{1})}$ associated with an admissible $R(\delta)$-perturbation on the space $\mathcal{L}^{\infty}(=\mathcal{L}^{\infty}_{m})$ is essential, as the measure $m$, which defines $\mathcal{L}^{\infty}$, is fixed. The $F_{\delta}$-invariant probabilities $\mu_{\delta}$ project onto $m_{\delta}$, which is absolutely continuous with respect to a conformal measure $\nu_{\delta}$ with Jacobian $\lambda_{\delta}e^{-\varphi}$. Moreover, the measure $m_{\delta}=h_{\delta}\nu_{\delta}$. The next result shows that a statement similar to Proposition 4.6 still holds uniformly for $\mathop{\mathrm{\overline{}}{F}}_{\Phi,\delta}$ for all $\delta$. Additionally, we recall that $\mathop{\mathrm{\overline{}}{F}}_{\Phi,\delta}$ is defined by Definition 4.3, using the abbreviated notation introduced in Remark 4.5.
@@
Suppose that $\{F_{\delta}\}_{\delta\in[0,1)}$ is an admissible $R(\delta)$-perturbation, and let $\mathop{\mathrm{\overline{}}{F}}_{\Phi,\delta}$ and $\mu_{\delta}$ denote the corresponding transfer operators and fixed points (i.e., the $F_{\delta}$ invariant probability measures in $S^{\infty}$), respectively. If the family $\{\mu_{\delta}\}_{\delta\in[0,1)}$ satisfies
@@
Consider an admissible $R(\delta)$-perturbation $\{F_{\delta}\}_{\delta\in[0,1)}$ and let $\mu_{\delta}$ be the family of $F_{\delta}$-invariant probability measures in $S^{\infty}$ (which is unique for each $\delta$), for all $\delta\in[0,1)$. Then, there exists $B_{u}>0$ such that
@@
According to Theorem 4.15, for each $\delta\in[0,1)$ we denote by $\mu_{\delta}\in S^{\infty}$ the unique $F_{\delta}$-invariant probability measure in $S^{\infty}$. For a fixed $\delta\in[0,1)$and following Definition 3.5, we consider a path
@@
By Theorem 4.17, the sequence $\{\mathop{\mathrm{\overline{}}{F}^{n}_{\Phi,\delta}}m_{1}\}_{n\in\mathbb{N}}$ converges to $\mu_{\delta}$ in $\mathcal{L}^{\infty}$, which implies that the sequence $\{\Gamma_{\delta}^{n}\}_{n\in\mathbb{N}}$ converges $m$-a.e. to $\Gamma_{\delta}\in\Gamma_{\mu_{\delta}}$ in $\mathcal{SB}(K)$ with respect to the metric defined in Definition 17. In particular, $\{\Gamma_{\delta}^{n}\}_{n\in\mathbb{N}}$ converges pointwise to $\Gamma_{\delta}$ on a full-measure subset $\widehat{M_{\delta}}\subset M$.
@@ 6.2. Perturbation of Operators
[…]
@@
Theorem D (Quantitative stability for deterministic perturbations).
Let $\{F_{\delta}\}_{\delta\in[0,1)}$ be an admissible $R(\delta)$-perturbation. Let $\mu_{0}$ be the unique equilibrium state of $\Phi\in\mathcal{P}$ in $S^{\infty}$. Denote by $\mu_{\delta}$ the invariant measure of $F_{\delta}$ in $S^{\infty}$, for all $\delta$. Then, there exist constants $D_{2}<0$ and $\delta_{1}\in(0,\delta_{0})$ such that for all $\delta\in[0,\delta_{1})$, it holds
@@
Let $\{F_{\delta}\}_{\delta\in[0,1)}$ be an admissible $R(\delta)$-perturbation, where $R(\delta)$ is defined by $R(\delta)=K_{5}\delta$. Denote by $\mu_{\delta}$ the unique invariant probability of $F_{\delta}$ in $S^{\infty}$, for all $\delta$. Then, there exist constants $D_{2}<0$ and $\delta_{1}\in(0,\delta_{0})$ such that for all $\delta\in[0,\delta_{1})$, it holds
@@ 7. Extending the Class of Potentials
A natural question is whether the results obtained so far can be extended to a broader class of potentials than $\mathcal{P}$. For instance, can they be extended to potentials that are not necessarily constant along the fibers? A possible approach is as follows.
Let $\mathcal{S}$ be a class of dynamical systems $F$ such that there exists a class of potentials $\overline{\mathcal{P}}$ with the property that, for every $\Phi\in\overline{\mathcal{P}}$, there exists a potential $\overline{\Phi}\in\mathcal{P}$ (see Equation (24)) for which the pairs $(F,\Phi)$ and $(F,\overline{\Phi})$ share the same equilibrium states.
In this case, all results established in the previous sections extend naturally to potentials in $\overline{\mathcal{P}}$. As we shall see, this phenomenon indeed occurs for an interesting class of dynamical systems and for an equally interesting class of potentials $\overline{\mathcal{P}}$.
As an immediate consequence, we obtain the following results.
Theorem E.
Suppose that $\in\mathcal{S}, FF$ satisfies (f1), (f2), (f3), (H1) and (H2), $(\alpha\cdot L)^{\zeta}<1$ and $\Phi\in\overline{\mathcal{P}}$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi$ in $S^{\infty}$. There exists a constant $0<\tau_{2}<1$ such that, for every constant fiber function $\psi:\Sigma\to\mathbb{R}$ satisfying $\psi(\cdot,y)\in L^{1}(m)$ for all $y$, and for every $\varphi\in\operatorname{H}_{\zeta}(\Sigma)$, we have
@@
Theorem F.
Suppose that $\in\mathcal{S}, FF$ satisfies (f1), (f2), (f3), (H1) and (H2), $(\alpha\cdot L)^{\zeta}<1$ and $\Phi\in\overline{\mathcal{P}}$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi$ in $S^{\infty}$. There exists a constant $0<\tau_{3}<1$ such that, for all $\psi\in L^{1}_{\mu_{0}}(\mathcal{F}_{0})$ and all $\varphi\in\operatorname{H}_{\zeta}(\Sigma)$, it holds
@@
Theorem G (Central Limit Theorem).
Suppose that $\in\mathcal{S}, FF$ satisfies (f1), (f2), (f3), (H1) and (H2), $(\alpha\cdot L)^{\zeta}<1$ and $\Phi\in\overline{\mathcal{P}}$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi$ in $S^{\infty}$. Given a Hölder continuous function $\varphi$, define
@@
Theorem H (Quantitative stability for deterministic perturbations).
Suppose that $\in\mathcal{S}, FF$ satisfies (f1), (f2), (f3), (H1) and (H2), $(\alpha\cdot L)^{\zeta}<1$ and $\Phi\in\overline{\mathcal{P}}$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi$ in $S^{\infty}$. Let $\{F_{\delta}\}_{\delta\in[0,1)}$ be an admissible $R(\delta)$-perturbation. Denote by $\mu_{\delta}$ the invariant measure of $F_{\delta}$ in $S^{\infty}$, for all $\delta$. Then, there exist constants $D_{2}<0$ and $\delta_{1}\in(0,\delta_{0})$ such that for all $\delta\in[0,\delta_{1})$, it holds
@@
7.1. An example of $\mathcal{S}$ and $\overline{\mathcal{P}}$
Define the family $\mathcal{S}$, by
$\mathcal{S}:=\{F:\Sigma\to\Sigma;F\ \text{is continuous},\ \exists\ y_{0}\in K\ \text{s.t.}\ G(x,y_{0})=y_{0}\ \forall\ x\in M\}.$ (53)

### v2

@@ 5.2. Exponential Decay of Correlations over $L^{1}(\mathcal{F}_{0})$
In this section, we introduce the space $L^{1}(\mathcal{F}_{0})$, on which we will later establish the exponential decay of correlations. This result will have a further application, where we prove the Central Limit Theorem for Hölder observables.
@@
Theorem C.
Suppose that $F:\Sigma\longrightarrow\Sigma$ satisfies (f1), (f2), (f3), (H1) and (H2) and $(\alpha\cdot L)^{\zeta}<1$. Let $\mu_{0}$ be the unique $F$-invariant probability in $\mathbf{S}^{\infty}$. There exists a constant $0<\tau_{3}<1$ such that, for all $\psi\in L^{1}_{\mu_{0}}(\mathcal{F}_{0})$ and all $\varphi\in\operatorname{H}_{\zeta}(\Sigma)$, it holds
@@
Note that if $\psi$ is a real-valued $\mathcal{F}_{0}$-measurable function, then $\psi$ is constant along the fibres. Indeed, $\psi^{-1}(x)$ is $\mathcal{F}_{0}$-measurable. If $\psi^{-1}(x)\neq\emptyset$, then there exist $\gamma\in\mathcal{F}^{s}$ and $z\in\gamma$ such that $\psi(z)=x$. By Proposition 5.3, we have $\gamma\subset\psi^{-1}(x)$, and therefore $\psi(z)=x$ for all $z\in\gamma$.
@@
By Theorem B, the proof is finished.
@@ 5.3. The Gordin’s Theorem
We now introduce an important tool that will be used, together with Theorem C, to establish the Central Limit Theorem: Gordin’s theorem. Its proof can be found in [21], and therefore we omit it here.
Let $(\Sigma,\mathcal{F}_{0},\mu)$ be a probability space where $\mathcal{F}_{0}$ is a $\sigma$-algebra of $\Sigma$. Let $F:\Sigma\longrightarrow\Sigma$ be a $\mathcal{F}_{0}$-measurable function and let $\mathcal{F}_{n}:=F^{-n}(\mathcal{F}_{0})$ be a non-increasing family of $\sigma$-algebras. A function $\xi:M\to\mathbb{R}$ is $\mathcal{F}_{n}$-measurable if and only if there exists a $\mathcal{F}$-measurable function $\xi_{n}$ such that $\xi=\xi_{n}\circ F^{n}$. Define
@@ 5.4. Application: Central Limit Theorem
By the Theorem C, if $\psi\in L^{1}_{\mu_{0}}(\mathcal{F}_{0})$ with $\int\psi\,d\mu\leq 1$, then
@@
Theorem D (Central Limit Theorem).
Suppose that $F:\Sigma\longrightarrow\Sigma$ satisfies (f1), (f2), (f3), (H1) and (H2) and $(\alpha\cdot L)^{\zeta}<1$. Let $\mu_{0}$ be the unique $F$-invariant probability in $\mathbf{S}^{\infty}$. Given a Hölder continuous function $\varphi$, define
@@ 6.1. Admissible $R(\delta)$-Perturbations
Since $\phi_{1}=1$ $m$-a.e., $\mu|_{\gamma_{\delta,i}}-\mu|_{\gamma}$ has zero average. Therefore, by lemmas 4.10 and 4.11, by (U2.2) and Definition (3.6) applied on $\mu$, we get
@@
for $m$-a.e. $\gamma\in M$ and all $\mu\in\mathbf{AB}_{m}$, where
@@
Understanding the action of the transfer operators $\{\mathop{\mathrm{\overline{}}{F}}_{\Phi,\delta}\}_{[0,\delta_{1})}$ associated with an admissible $R(\delta)$-perturbation on the space $\mathbf{L}^{\infty}(=\mathbf{L}^{\infty}_{m})$ is essential, as the measure $m$, which defines $\mathbf{L}^{\infty}$, is fixed. The $F_{\delta}$-invariant probabilities $\mu_{\delta}$ project onto $m_{\delta}$, which is absolutely continuous with respect to a conformal measure $\nu_{\delta}$ with Jacobian $\lambda_{\delta}e^{-\varphi}$. Moreover, the measure $m_{\delta}=h_{\delta}\nu_{\delta}$. The next result shows that a statement similar to Proposition 4.7 still holds uniformly for $\mathop{\mathrm{\overline{}}{F}}_{\Phi,\delta}$ for all $\delta$. Additionally, we recall that $\mathop{\mathrm{\overline{}}{F}}_{\Phi,\delta}$ is defined by Definition 4.4, using the abbreviated notation introduced in Remark 4.6.
@@
Suppose that $\{F_{\delta}\}_{\delta\in[0,1)}$ is an admissible $R(\delta)$-perturbation, and let $\mathop{\mathrm{\overline{}}{F}}_{\Phi,\delta}$ and $\mu_{\delta}$ denote the corresponding transfer operators and fixed points (i.e., the $F_{\delta}$ invariant probability measures in $\mathbf{S}^{\infty}$), respectively. If the family $\{\mu_{\delta}\}_{\delta\in[0,1)}$ satisfies
@@
Consider an admissible $R(\delta)$-perturbation $\{F_{\delta}\}_{\delta\in[0,1)}$ and let $\mu_{\delta}$ be the family of $F_{\delta}$-invariant probability measures in $\mathbf{S}^{\infty}$ (which is unique for each $\delta$), for all $\delta\in[0,1)$. Then, there exists $B_{u}>0$ such that
@@
According to Theorem 4.16, for each $\delta\in[0,1)$ we denote by $\mu_{\delta}\in\mathbf{S}^{\infty}$ the unique $F_{\delta}$-invariant probability measure in $\mathbf{S}^{\infty}$. For a fixed $\delta\in[0,1)$and following Definition 3.5, we consider a path
@@
By Theorem 4.17, the sequence $\{\mathop{\mathrm{\overline{}}{F}^{n}_{\Phi,\delta}}m_{1}\}_{n\in\mathbb{N}}$ converges to $\mu_{\delta}$ in $\mathbf{L}^{\infty}$, which implies that the sequence $\{\Gamma_{\delta}^{n}\}_{n\in\mathbb{N}}$ converges $m$-a.e. to $\Gamma_{\delta}\in\Gamma_{\mu_{\delta}}$ in $\mathcal{SB}(K)$ with respect to the metric defined in Definition 17. In particular, $\{\Gamma_{\delta}^{n}\}_{n\in\mathbb{N}}$ converges pointwise to $\Gamma_{\delta}$ on a full-measure subset $\widehat{M_{\delta}}\subset M$.
@@ 6.2. Perturbation of Operators
Let $\{F_{\delta}\}_{\delta\in[0,1)}$ be an admissible $R(\delta)$-perturbation and let $\{\overline{\mathop{\mathrm{F}}}_{\Phi,\delta}\}_{\delta\in[0,1)}$ be the induced family of transfer operators. Then, $\{\overline{\mathop{\mathrm{F}}}_{\Phi,\delta}\}_{\delta\in[0,1)}$ is an $(R(\delta),\zeta)$-family of operators with $(\mathbf{L}^{\infty},||\cdot||_{\infty})$ and $(\mathbf{S}^{\infty},||\cdot||_{\mathbf{S}^{\infty}})$ as the weak and strong spaces of the family, respectively.
@@
Theorem E (Quantitative stability for deterministic perturbations).
Let $\{F_{\delta}\}_{\delta\in[0,1)}$ be an admissible $R(\delta)$-perturbation. Let $\mu_{0}$ be the unique $F$-invariant probability in $\mathbf{S}^{\infty}$. Denote by $\mu_{\delta}$ the invariant measure of $F_{\delta}$ in $\mathbf{S}^{\infty}$, for all $\delta$. Then, there exist constants $D_{2}<0$ and $\delta_{1}\in(0,\delta_{0})$ such that for all $\delta\in[0,\delta_{1})$, it holds
@@
Let $\{F_{\delta}\}_{\delta\in[0,1)}$ be an admissible $R(\delta)$-perturbation, where $R(\delta)$ is defined by $R(\delta)=K_{5}\delta$. Denote by $\mu_{\delta}$ the unique invariant probability of $F_{\delta}$ in $\mathbf{S}^{\infty}$, for all $\delta$. Then, there exist constants $D_{2}<0$ and $\delta_{1}\in(0,\delta_{0})$ such that for all $\delta\in[0,\delta_{1})$, it holds
@@ 7. Extending the Class of Potentials
A natural question is whether the results obtained so far extend to a broader class of potentials than $\mathscr{P}_{\Sigma}$ (see Equation (24)), in particular to potentials that are not necessarily constant along the fibres. To address this question, and to isolate it as a co-homological problem, we introduce the class $\mathcal{S}$ of dynamical systems $F$ satisfying the following properties:
  1. (1)
    $F$ satisfies assumptions (f1), (f2), and (f3), and uniformly contracts all vertical fibres;
  2. (2)
    there exists a class of potentials $\overline{\mathscr{P}}_{\Sigma}$ such that, for every $\overline{\Phi}\in\overline{\mathscr{P}}_{\Sigma}$, one can associate a potential $\Phi\in\mathscr{P}_{\Sigma}$ for which the pairs $(F,\Phi)$ and $(F,\overline{\Phi})$ share the same equilibrium states.
Note that, in order to belong to the class $\mathcal{S}$, the map $G$ must satisfy a condition stronger than (H1).
In this case, all results established in the previous sections extend naturally to potentials in $\overline{\mathscr{P}}_{\Sigma}$. As we shall see, this phenomenon indeed occurs for an interesting class of dynamical systems and for an equally interesting class of potentials $\overline{\mathscr{P}}_{\Sigma}$.
As an immediate consequence, we obtain the following results. Note that, by the definition of $\mathcal{S}$ and Theorem 4.16, in Theorems E–H the probability measure $\mu_{0}$ is an equilibrium state for the potential $\Phi\in\mathscr{P}_{\Sigma}$.
Theorem F.
Suppose that $\in\mathcal{S}, FF$ satisfies (f1), (f2), (f3), (H2), $(\alpha\cdot L)^{\zeta}<1$ and $\Phi\in\overline{\mathscr{P}}_{\Sigma}$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi$ in $\mathbf{S}^{\infty}$. There exists a constant $0<\tau_{2}<1$ such that, for every constant fiber function $\psi:\Sigma\to\mathbb{R}$ satisfying $\psi(\cdot,y)\in L^{1}(m)$ for all $y$, and for every $\varphi\in\operatorname{H}_{\zeta}(\Sigma)$, we have
@@
Theorem G.
Suppose that $\in\mathcal{S}, FF$ satisfies (f1), (f2), (f3), (H2), $(\alpha\cdot L)^{\zeta}<1$ and $\Phi\in\overline{\mathscr{P}}_{\Sigma}$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi$ in $\mathbf{S}^{\infty}$. There exists a constant $0<\tau_{3}<1$ such that, for all $\psi\in L^{1}_{\mu_{0}}(\mathcal{F}_{0})$ and all $\varphi\in\operatorname{H}_{\zeta}(\Sigma)$, it holds
@@
Theorem H (Central Limit Theorem).
Suppose that $\in\mathcal{S}, FF$ satisfies (f1), (f2), (f3), (H2), $(\alpha\cdot L)^{\zeta}<1$ and $\Phi\in\overline{\mathscr{P}}_{\Sigma}$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi$ in $\mathbf{S}^{\infty}$. Given a Hölder continuous function $\varphi$, define
@@
Theorem I (Quantitative stability for deterministic perturbations).
Suppose that $\in\mathcal{S}, FF$ satisfies (f1), (f2), (f3), (H2), $(\alpha\cdot L)^{\zeta}<1$ and $\Phi\in\overline{\mathscr{P}}_{\Sigma}$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi$ in $\mathbf{S}^{\infty}$. Let $\{F_{\delta}\}_{\delta\in[0,1)}$ be an admissible $R(\delta)$-perturbation. Denote by $\mu_{\delta}$ the invariant measure of $F_{\delta}$ in $\mathbf{S}^{\infty}$, for all $\delta$. Then, there exist constants $D_{2}<0$ and $\delta_{1}\in(0,\delta_{0})$ such that for all $\delta\in[0,\delta_{1})$, it holds
@@
7.1. An example of $\mathcal{S}$ and $\overline{\mathscr{P}}_{\Sigma}$
In what follows, $\operatorname{H}_{\zeta}(\Sigma)$ denotes the space of real-valued $\zeta$-Hölder continuous functions defined on $\Sigma:=M\times K$. Throughout this discussion, we assume $F$ to be continuous.
For any $\bar{\Phi}\in\operatorname{H}_{\zeta}(\Sigma)$ and a fixed element $y_{0}\in K$ (as specified in Equation (53)), we define the function $\bar{\Phi}_{y_{0}}:M\to\mathbb{R}$ by $\bar{\Phi}_{y_{0}}(x):=\bar{\Phi}(x,y_{0})$. We then introduce the family $\mathcal{S}$ of maps as follows:
$\mathcal{S}:=\{F:\Sigma\to\Sigma\mid F\text{ is continuous and }\exists\,y_{0}\in K\text{ s.t. }G(x,y_{0})=y_{0},\,\forall\,x\in M\}.$ (53)
Furthermore, we define the set of potentials $\overline{\mathscr{P}}_{\Sigma}\subset\operatorname{H}_{\zeta}(\Sigma)$ by
$\overline{\mathscr{P}}_{\Sigma}:=\{\bar{\Phi}\in\operatorname{H}_{\zeta}(\Sigma)\mid\bar{\Phi}_{y_{0}}\in\mathscr{P}_{M}\}.$ (54)